A company offers a raffle whose grand prize is a $50,000 new car. Additional prizes are a $1,000 television and a $500 computer. Tickets cost $20 each. Ticket income will be donated to a charity. If 3,000 tickets are sold, what is the expected gain or loss (in dollars) of each ticket? (Round your answer to the nearest cent.)

I would appreciate it if the answer was written electronically.
Answer #1

Let X be the prize amount. Total no of tickets sold = 3000

P(X = $ 50000) = 1/3000 , P(X = $1000) = 1/3000 , P(X = $ 500) = 1/3000

Expected prize amount = xP(X = x) = 50000/3000 + 1000/3000 + 500/3000 = $ 17.17

Cost of each ticket = $ 20

So, Expected gain/Loss = 17.17 - 20.00 = $ -2.83 (Negative sign shows a loss)